# An Empirical Evaluation of Pre-trained Large Language Models for Repairing Declarative Formal Specifications

### v1

@@ 3.2.1. Repair Agent
3.2.2. Prompt Agent
This agent has its LLM settings and context and generates feedback based on the report generated by Alloy Analyzer and the proposed specification. The produced feedback guides the Repair Agent on the methods for rectifying the bug. The system message that provides the instructions to this agent is described in Table 1. Unlike the system message of the Repair Agent, the agent’s persona is described as “Expert in Analyzing Alloy Analyzer reports”. This agent does not share the same context or prompts with the Repair Agent. But the response of the Prompt Agent is appended to the prompt that will be used in the next iteration by the Repair Agent.
3.3. Feedback Messages
The initial prompt contains only Agent Instruction and Tool Instruction defined by the Repair Agent, as described in Table 1. The pipeline refines the prompts in response to the behavior of the proposed specification. This refinement process is contingent upon repetitive buggy specifications, errors that emerge after the execution of these specifications through the Alloy analyzer, or LLM failure to send a response based on the required JSON format.
@@ 3.3. Feedback Messages
  • Tool fallback. This message informs LLM that the response received does not comply with the tool’s JSON format, consequently impeding the extraction of the proposed specification. To address this issue, we developed a parser atop the Langroid JSON parser, specifically to retrieve Alloy specifications from the response. Nonetheless, the parser cannot handle all scenarios, resulting in the dispatch of this feedback to the LLM, soliciting adherence to the mandated JSON format.
@@
Table 2. Feedback Messages used by the repair pipeline
@@
Tool-fallback Alloy Analyzer You must use the CORRECT format described in the tool ‘run_alloy_analyzer‘ to send me the fixed specifications. You either forgot to use it, or you used it with the WRONG format. Make sure all fields are filled out.
Repeated spec The proposed ¡FIXED_SPECIFICATIONS¿ is IDENTICAL to the Alloy ¡Faulty_SPECIFICATIONS¿ that I sent you. **DO NOT** send Alloy specifications that I sent you again. ALWAYS USE the tool ‘run_alloy_analyzer‘ to send me a new ¡FIXED_SPECIFICATIONS¿.
@@
Generic-feedback Below are the results from the Alloy Analyzer. Fix all Errors and Counterexamples before sending me the next ¡FIXED_SPECIFICATIONS¿.
Auto-feedback Prompt Agent The message content will be dynamically constructed by the agent based on the supplied information.
@@
  3. (3)
    Auto-feedback. When this setting is enabled, the feedback response to LLM will be generated by another LLM agent (i.e., prompt agent). We forward the report generated by the Alloy Analyzer and the proposed Alloy specifications to the prompt agent, which generates a prompt to instruct the Repair Agent.
The different reports reflect the ability of LLMs to (1) locate bugs and (2) correctly perform the repair process. For example, Generic-feedback and Auto-feedback show the ability of LLM to perform its reasoning based on a human-created prompt versus an LLM-generated prompt, which has no access to the repair context. Also, the latter reflects a real scenario wherein the user cannot repair the buggy specification, then they consult experts or other forums to get assistance.
4. Experiment Design
@@
  • RQ1 (Effectiveness): How effective is the APR pipeline in repairing compared to the state of the art?
@@
  • RQ3 (Adaptive Prompting): What is the impact of applying the adaptive prompting technique?
@@
  • RQ4 (Failure Characteristics): What are the characteristics of failures encountered during the repair process?
@@
  • RQ5 (Repair Costs): What is the cost associated with using the APR pipeline with various LLMs?
4.1. Dataset
In our evaluation, we utilize two distinct benchmark suites: ARepair (Wang et al., 2019) and Alloy4Fun (Macedo et al., 2021). These benchmark suites have undergone extensive study and been developed by independent research groups, enabling a fair comparison of various techniques.
@@ 4.1. Dataset
The benchmark datasets consist of specifications varying from tens to hundreds of lines, featuring real bugs authored by humans. The defects within the benchmarks span a diverse range, from straightforward issues that can be addressed by modifying a single operator to complex challenges necessitating the synthesis of novel expressions and the substitution of entire predicate bodies.
@@
The ARepair benchmark (Wang et al., 2019) encompasses 38 faulty specifications extracted from a total of 12 Alloy problems. Within this benchmark, both single and multi-line errors are present, distributed across 28 faulty models with single-line errors and 10 faulty models with multi-line errors. The Alloy4Fun benchmark (Macedo et al., 2021) comprises a collection of 1,938 handwritten defective models sourced from student submissions across six different Alloy problems. All bugs within this benchmark are single-line bugs. Both benchmarks are accompanied by correct versions of the specifications, serving as ground truths for verifying the accuracy of the obtained results.
4.2. Pre-processing Benchmark Dataset
@@
As mentioned earlier, the benchmark datasets include comments indicating bug locations and their corresponding fixes. To ensure a fair evaluation, we have removed these comments from the defective models. Specifically, we have eliminated all lines starting with the phrase ”Fix:” in the defective models. This line typically outlines the necessary changes required to resolve the bug. Models with multiple ”Fix:” comments are categorized as having multi-line bugs, whereas those with only one ”Fix:” comment are considered to have single-line bugs.
This process of removing comments is crucial as it prevents LLMs from being provided with explicit clues about the locations of bugs and their respective fixes. By doing so, we aim to accurately assess the repair capabilities of LLMs and observe their effectiveness in localizing bugs within the defective models.
4.3. Implementation
@@
Our implementation of the APR pipeline was realized using Python 3.11, with the assistance of Langroid to facilitate the integration of multi-agent LLMs. This implementation encompasses several crucial functionalities. Firstly, it offers support for various LLMs, including local LLMs, ensuring flexibility in model selection. Secondly, it incorporates a message history control mechanism, essential for preventing context length limitations, particularly during iterative repair processes. Lastly, our implementation facilitates the creation and customization of tools, allowing for the definition of response formats and fields. Additionally, we developed a specialized parser to address issues caused by special situations that could impede the extraction of proposed specifications.
@@
4.4. Subject LLMs
To assess the performance of existing pre-trained LLMs, we have carefully chosen three representative models [^5] While exploring local models such as Llama-2 and Mistral, it became apparent that they lacked support for tools, making it challenging to process their responses, which often contained incomplete Alloy specifications.:
  • GPT-3.5-Turbo: This standard LLM, provided by OpenAI and utilized in ChatGPT, boasts a sophisticated architecture with 175 billion parameters, endowing it with extensive capabilities. Engineered to tackle a wide array of natural language processing (NLP) tasks, including text generation and completion, GPT-3.5-Turbo epitomizes advanced computational linguistics. Within the context of our study, we utilize the gpt-3.5-Turbo variant accessed via the OpenAI API for evaluative purposes.
  • GPT-4-32k: Representing the state-of-the-art in LLM models from OpenAI, this model contains over 1.76 trillion parameters, demonstrating significantly enhanced capabilities compared to GPT-3.5. For our study, we leverage the GPT-4-32k-0613 model version.
  • GPT-4-Turbo: Another iteration of the GPT-4 model, GPT-4-Turbo features an updated knowledge cutoff as of April 2023 and introduces a 128k context window. Moreover, it offers cost-effectiveness compared to GPT-4, alongside notable enhancements such as improved instruction following, JSON mode, and reproducible outputs (OpenAI, [n. d.]).
Table 3 provides more details about the characteristics of these pre-trained LLMs.
Table 3. Characteristics of Pre-trained LLMs
@@ 4.4. Subject LLMs
4.5. Subject SOTA Systems
@@
We conducted a comparative analysis of repair performance against several state-of-the-art Alloy repair tools: ARepair (Wang et al., 2019), ICEBAR (Gutiérrez Brida et al., 2023), BeAFix (Brida et al., 2021), ATR (Zheng et al., 2022), and Hasan et al. (Hasan et al., 2023). Each tool employs a distinct development approach and tackles specification attributes differently.
@@
ARepair (Wang et al., 2019) generates fixes for Alloy specifications that violate test cases. ICEBAR (Gutiérrez Brida et al., 2023) utilizes ARepair as a backend tool to repair faulty Alloy specifications based on a predefined set of Alloy tests. ATR (Zheng et al., 2022) adopts a template-based methodology to enhance the repair process, leveraging fault localization strategies to identify potentially erroneous Alloy expressions leading to assertion failures. BeAFix (Brida et al., 2021) relies on user input to identify faulty statements. It exhaustively explores all possible repair candidates up to a certain threshold through mutation and employs Alloy counterexamples to evaluate the feasibility of generalization. Fianlly, Hasan et al. (Hasan et al., 2023) employed ChatGPT for repairing faulty specifications across five scenarios. For our comparison, we exclude scenarios where ChatGPT is provided with “Fix:” comments, as this represents an unrealistic setting and does not align with our ”Benchmark Pre-processing” step (cf. Section 4.2).
4.6. Experimental Configuration and Settings
@@
LLM Temperature: To balance between deterministic progression and iterative refinement, we set the temperature parameter of the Large Language Models (LLMs) to 0.2, allowing for a moderate level of randomness while maintaining some level of determinism throughout the repair process (Kong et al., 2024; Zhang et al., 2024).
@@
Number of Iterations: Initially, we conducted a preliminary experiment using the ARepair benchmark framework with GPT-4-32k, allocating a budget of ten iterations. However, we observed diminishing returns after six iterations in most cases. Therefore, we opted for a six-iteration budget in the APR pipeline.
@@
Metric for Comparing Repair Performance of LLMs: We employ the Correct@k metric (Liu et al., 2024) to evaluate the success rate of the techniques in repairing defective Alloy specifications. This metric quantifies the number of defects successfully repaired within a maximum of $k$ iterations, where $k$ is set to 6 in our study (denoted as Correct@6). The formula for this metric is described in Equation 1.
@@
Experiment Settings and Baseline: Our evaluation encompasses a total of 1,974 defective models. We assess three distinct LLMs, as outlined in Table 3. The APR pipeline supports three levels of report granularity and permits multiple iterations, resulting in a total of 106,596 repair attempts (calculated as the product of 1,974 benchmarks, three LLMs, three feedback levels, and up to six iterations). To manage costs associated with GPT-4 models, we conducted two experiments for RQ1. The first experiment focuses on repairing Alloy4Fun and ARepair benchmarks using GPT-3.5, followed by a comparison with state-of-the-art (SoTA) tools. In the second experiment, we limit the repair task to the ARepair benchmark, utilizing GPT-4 and GPT-4-Turbo exclusively. Table 4 provides details of the different combinations of LLMs and feedback intensities used in the latter experiment.
Table 4. Settings corresponding to various combinations of LLMs and feedback levels
Setting No. LLM Feedback Level
Setting-1 GPT-4-32k No-Feedback
Setting-2 GPT-4-32k Generic-Feedback
Setting-3 GPT-4-32k Auto-Feedback
Setting-4 GPT-4-Turbo No-Feedback
Setting-5 GPT-4-Turbo Generic-Feedback
Setting-6 GPT-4-Turbo Auto-Feedback
5. Experimental Results

### v2

@@ 3.2.1 Repair Agent
3.2.2 Instructor Agent
This agent has its LLM settings and context and generates feedback based on the report generated by Alloy Analyzer and the proposed specification. This agent leverages the AutoPrompt concept Shin et al. (2020) by automatically constructing a prompt using LLMs. The produced feedback (i.e., the automatically constructed prompt) guides the Repair Agent on the methods for rectifying the bug. The system message that provides the instructions to this agent is described in Table 1. Unlike the system message of the Repair Agent, the agent’s persona is described as “Expert in Analyzing Alloy Analyzer reports”. This agent does not share the same context or prompts with the Repair Agent. But the response of the Instructor Agent is appended to the prompt that will be used in the next iteration by the Repair Agent, as illustrated in Figure 1.
3.3 Feedback Messages
The initial prompt contains only Agent Instruction and Tool Instruction defined by the Repair Agent, as described in Table 1. The pipeline refines the prompts in response to the behavior of the proposed specification. This refinement process is contingent upon repetitive buggy specifications, errors that emerge after executing these specifications through the Alloy analyzer, or LLM failure to send a response based on the required JSON format.
@@ 3.3 Feedback Messages
  • Tool fallback. This message informs LLM that the response received does not comply with the tool’s JSON format, impeding the proposed specification’s extraction. To address this issue, we developed a parser atop the Langroid JSON parser to retrieve Alloy specifications from the response. Nonetheless, the parser cannot handle all scenarios, resulting in the dispatch of this feedback to the LLM, soliciting adherence to the mandated JSON format.
@@
Table 2: Feedback Messages used by the repair pipeline
@@
Tool-fallback Repair Agent You must use the CORRECT format described in the tool ‘run_alloy_analyzer‘ to send me the fixed specifications. You either forgot to use it, or you used it with the WRONG format. Make sure all fields are filled out.
Repeated spec The proposed $<$FIXED_SPECIFICATIONS$>$ is IDENTICAL to the Alloy $<$Faulty_SPECIFICATIONS$>$ that I sent you. **DO NOT** send Alloy specifications that I sent you again. ALWAYS USE the tool ‘run_alloy_analyzer‘ to send me a new $<$FIXED_SPECIFICATIONS$>$.
@@
Generic-feedback Below are the results from the Alloy Analyzer. Fix all Errors and Counterexamples before sending me the next $<$FIXED_SPECIFICATIONS$>$.
Auto-feedback Instructor Agent The message content will be dynamically constructed by the agent based on the supplied information.
@@
  3. 3.
    Auto-feedback. When this setting is enabled, the feedback response to LLM will be generated by another LLM agent (i.e., Instructor agent), which represents the dual-agent repair pipeline. We forward the report generated by the Alloy Analyzer and the proposed Alloy specifications to the prompt agent, which generates a prompt to instruct the Repair Agent.
Noteworthy, the No-feedback and Generic-feedback mechanisms are used under the single-agent paradigm, whereas the Auto-feedback approach enables the dual-agent paradigm. The different reports reflect the ability of LLMs to (1) locate bugs and (2) correctly perform the repair process. For example, Generic-feedback and Auto-feedback show the ability of LLM to perform its reasoning based on a human-created prompt versus an LLM-generated prompt, which has no access to the repair context. Also, the latter reflects a real scenario wherein the user cannot repair the buggy specification, then they consult experts or other forums to get assistance.
4 Experiment Design
@@
  • RQ1 (Effectiveness): How effective is the APR pipeline in repairing compared to the state-of-the-art and how various repair settings contribute to the effectiveness?
@@
  • RQ3 (Failure Characteristics): What are the characteristics of failures encountered during the repair process?
@@
  • RQ4 (Repair Costs): What is the cost associated with using the APR pipeline with various LLMs?
4.1 Dataset
@@
In our evaluation, we utilize two distinct benchmark suites: ARepair Wang et al. (2019) and Alloy4Fun Macedo et al. (2021). These benchmark suites have undergone extensive study and developed by independent research groups, enabling a fair comparison of various techniques. The benchmark datasets consist of specifications varying from tens to hundreds of lines, featuring real bugs authored by humans. The defects within the benchmarks span a diverse range, from straightforward issues that can be addressed by modifying a single operator to complex challenges necessitating the synthesis of novel expressions and the substitution of entire predicate bodies.
@@
The ARepair benchmark Wang et al. (2019) encompasses 38 faulty specifications extracted from a total of 12 Alloy problems. Within this benchmark, both single and multi-line errors are present, distributed across 28 faulty models with single-line errors and 10 faulty models with multi-line errors. The Alloy4Fun benchmark Macedo et al. (2021) comprises a collection of 1,938 handwritten defective models sourced from student submissions across six different Alloy problems. All bugs within this benchmark are single-line bugs. Both benchmarks are accompanied by correct versions of the specifications, serving as ground truths for verifying the accuracy of the obtained results.
4.2 Pre-processing Benchmark Dataset
As mentioned earlier, the benchmark datasets include comments indicating the locations of the bugs and their corresponding fixes. To ensure a fair evaluation, we perform two steps: (1) determining the uniqueness of Alloy specifications and (2) removing fix comments from the defective models.
Uniqueness of Alloy Specifications. To ensure that all specifications in the ARepair and Alloy4Fun benchmarks are unique, we applied a systematic normalization and hashing process. Each specification was processed in its entirety, preserving the original content. We removed block comments, and collapsed all whitespace (spaces, tabs, line breaks) to a single space. This normalization step eliminates superficial differences such as formatting and comments, allowing us to focus on the core logic of each Alloy specification. After normalization, we computed an MD5 hash for each specification, treating the hash as a unique fingerprint. Specifications with identical hashes were considered duplicates. We grouped specifications by hash and maintained counts for each benchmark and for the combined dataset. Our analysis found that all 38 ARepair specifications and all 1936 Alloy4Fun files were unique—no duplicates were detected in the combined set of 1974 specifications. This confirms the diversity and uniqueness of the Alloy specifications in both datasets.
@@ 4.2 Pre-processing Benchmark Dataset
Removal of Fix comments. We have removed these comments from the defective models. Specifically, we have eliminated all lines starting with the phrase “Fix:” in the defective models. This line typically outlines the necessary changes required to resolve the bug. Models with multiple “Fix:” comments are categorized as having multi-line bugs, whereas those with only one “Fix:” comment are considered to have single-line bugs. Removing comments is essential to prevent LLMs from receiving explicit clues about bug locations and fixes. This ensures an accurate assessment of their repair capabilities and effectiveness in localizing bugs within defective models. Importantly, this reflects a realistic scenario in which the user employs LLMs without prior knowledge of the bug location and the fix.
4.3 Implementation
@@
Our implementation of the APR pipeline was realized using Python 3.11, with the assistance of Langroid to facilitate the integration of multi-agent LLMs. This implementation encompasses several crucial functionalities. Firstly, it offers support for various LLMs, including local LLMs, ensuring flexibility in model selection. Secondly, it incorporates a message history control mechanism, essential for preventing context-length limitations, particularly during iterative repair processes. Lastly, our implementation facilitates the creation and customization of tools, allowing for the definition of response formats and fields. Additionally, we developed a specialized parser to address issues caused by special situations that could impede the extraction of proposed specifications.
@@
4.4 Subject LLMs
To assess the performance of existing pre-trained LLMs, we have carefully chosen three representative models. We explored local LLMs such as Llama-2 and Mistral, but they lacked support for tools and JSON responses, making it challenging to process their responses, which often contained incomplete Alloy specifications. Following is a description of the selected models, which are summarized in Table 3:
  • GPT-3.5-Turbo. This standard LLM, provided by OpenAI and utilized in ChatGPT, boasts a sophisticated architecture with 175 billion parameters, endowing it with extensive capabilities. Engineered to tackle a wide array of natural language processing (NLP) tasks, including text generation and completion, GPT-3.5-Turbo epitomizes advanced computational linguistics.
  • GPT-4-32k. This model contains over 1.76 trillion parameters, demonstrating significantly enhanced capabilities compared to GPT-3.5. For our study, we leverage the GPT-4-32k-0613 model version.
  • GPT-4-Turbo. Another iteration of the GPT-4 model, GPT-4-Turbo features an updated knowledge cutoff as of April 2023 and introduces a 128k context window. Moreover, it offers cost-effectiveness compared to GPT-4, alongside notable enhancements such as improved instruction following, JSON mode, and reproducible outputs OpenAI (2023a).
  • GPT-4o. This model shares features with GPT-4-Turbo but is more cost-effective and has an updated knowledge base as of October 2023.
Table 3: Characteristics of Pre-trained LLMs
@@ 4.4 Subject LLMs
GPT-4o 2024-05-01-preview Oct 2023 128k \$5 \$15
4.5 Subject SOTA Systems
@@
We conducted a comparative analysis of repair performance against several state-of-the-art Alloy repair tools: ARepair Wang et al. (2019), ICEBAR Gutiérrez Brida et al. (2023), BeAFix Brida et al. (2021), ATR Zheng et al. (2022), and Hasan et al. Hasan et al. (2023). Each tool employs a distinct development approach and tackles specification attributes differently.
@@
ARepair Wang et al. (2019) generates fixes for Alloy specifications that violate test cases. ICEBAR Gutiérrez Brida et al. (2023) utilizes ARepair as a backend tool to repair faulty Alloy specifications based on a predefined set of Alloy tests. ATR Zheng et al. (2022) adopts a template-based methodology to enhance the repair process, leveraging fault localization strategies to identify potentially erroneous Alloy expressions that lead to assertion failures. BeAFix Brida et al. (2021) relies on user input to identify faulty statements. It exhaustively explores all possible repair candidates up to a certain threshold through mutation and employs Alloy counterexamples to evaluate the feasibility of generalization. Finally, Hasan et al. Hasan et al. (2023) employed ChatGPT (GPT-3.5-Turbo) for repairing faulty specifications across five scenarios. This approach employs a purely LLM-based method, wherein a single repair iteration is conducted without the implementation of agent setup or feedback mechanisms. For our comparison, we exclude scenarios where ChatGPT is provided with “Fix:” comments, as this represents an unrealistic setting and does not align with our “Benchmark Pre-processing” step (cf. Section 4.2).
Since the state-of-the-art tools—ARepair, ICEBAR, BeAFix, and ATR—were evaluated on static Alloy models prior to the release of Alloy 6, our experimental analysis focused exclusively on static Alloy models to ensure a fair comparison. These tools do not support mutable specifications or recently introduced features such as linear temporal logic (LTL), making them incompatible with models that leverage Alloy 6’s new capabilities. Additionally, due to structural dependencies and the absence of updated implementations, these tools could not be reliably executed on datasets incorporating Alloy 6-specific features. Nevertheless, our repair technique is not inherently limited to static specifications and, in principle, can be extended to support the dynamic capabilities introduced in Alloy 6.
4.6 Experimental Configuration and Settings
@@
LLM Temperature. To balance between deterministic progression and iterative refinement, we set the temperature parameter of the Large Language Models (LLMs) to 0.2, allowing for a moderate level of randomness while maintaining some level of determinism throughout the repair process Kong et al. (2024); Zhang et al. (2024).
@@
Number of Iterations. Initially, we conducted a preliminary experiment using the ARepair benchmark framework with GPT-4-32k, allocating a budget of ten iterations. However, we observed diminishing returns after six iterations in most cases. Therefore, we opted for a six-iteration budget in the APR pipeline.
@@
Metric for Comparing Repair Performance of LLMs. We employ the Correct@k metric Liu et al. (2024) to evaluate the success rate of the techniques in repairing defective Alloy specifications. This metric quantifies the number of defects successfully repaired within a maximum of $k$ iterations, where $k$ is set to 6 in our study (denoted as Correct@6). The formula for this metric is described in Equation 1.
@@
Experiment Settings and Baseline. Our evaluation encompasses a total of 1,974 defective models. We assess three distinct LLMs, as outlined in Table 3. The APR pipeline supports three levels of report granularity and permits multiple iterations, resulting in a total of 106,596 repair attempts (calculated as the product of 1,974 benchmarks, three LLMs, three feedback levels, and up to six iterations). The Single-agent paradigm utilizes the No-feedback and Generic- feedback mechanisms, whereas the dual-agent paradigm employs the Auto-feedback approach within the repair pipeline.
The repair experiment on the Alloy4Fun benchmark leverages GPT-4o—the most cost-effective option among GPT-4-32k and GPT-4-Turbo—for a randomly sampled subset of 357 models, maximizing both efficiency and performance. For comprehensive coverage, GPT-3.5-Turbo is applied to the entire Alloy4Fun dataset, taking advantage of its substantially lower cost compared to the GPT-4 family. This experimental design enables a broad and cost-efficient evaluation across all models, while also allowing for targeted, high-performance analysis using GPT-4o on a representative subset. Table 4 presents the settings corresponding to different combinations of LLMs and feedback levels.
Table 4: Settings corresponding to various combinations of LLMs and feedback levels
Setting No. # Agents LLM Feedback Level
Setting-1 Single GPT-4-32k No-Feedback
Setting-2 Single GPT-4-32k Generic-Feedback
Setting-3 Dual GPT-4-32k Auto-Feedback
Setting-4 Single GPT-4-Turbo No-Feedback
Setting-5 Single GPT-4-Turbo Generic-Feedback
Setting-6 Dual GPT-4-Turbo Auto-Feedback
Setting-7 Single GPT-4o No-Feedback
Setting-8 Single GPT-4o Generic-Feedback
Setting-9 Dual GPT-4o Auto-Feedback
Setting-10 Single GPT-3.5-Turbo No-Feedback
Setting-11 Single GPT-3.5-Turbo Generic-Feedback
Setting-12 Dual GPT-3.5-Turbo Auto-Feedback
5 Experimental Results